OpenStudy (anonymous):

Find the root. For g(t), find two roots nearest x=0. For h(x), find both roots. F(x)=-5x+3 and g(t)=1.707cos(2πt) and h(x)= 2x^2+x-12

OpenStudy (mathmale):

would you please tell us what methods you've been using to do similar problems. I suspect you're learning Newton's Method, which is a great approach. Once I've heard back from you, I'd be glad to proceed further towards helping you solve these problems.

OpenStudy (anonymous):

we are using mathcad it programms it for us i just need to know how to set it up

OpenStudy (anonymous):

ermm.. i don't kno really :-/

OpenStudy (mathmale):

If you're open to applying Newton's Method to finding the roots of the 2nd 2 equations, I can help. The first equation is just a straight line; to find the roo of that, just let x=0 and solve the resulting equation for x.

OpenStudy (mathmale):

Please set F(x)=-5x+3 equal to zero. Drop the label F(x). Solve the resulting equation for x. that value is the root you wanted.

OpenStudy (anonymous):

-5x+3=0. so x will be equal to3/5

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

that's right. The "root" or "zero" of the first equation is x=3/5. Now go on to the next problem. "for g(t), find two roots nearest x=0, if g(t)=1.707cos(2πt)

OpenStudy (mathmale):

I strongly advise you to use Newton's Method here. The work is a lot more challenging than it was for the first equation. Are you familiar with Newton's Method?

OpenStudy (anonymous):

no i'm not familar with newtons law

OpenStudy (mathmale):

Actually, it's Newton's Method, not Newton's 1st, 2nd and 3rd Laws (of Physics). The third problem is a bit easier to solve than is the 2nd. h(x)= 2x^2+x-12 could be set equal to zero and solved by either factoring or through the quadratic formula. If 2x^2+x-12=0, how would you go about finding the two roots/zeros?

OpenStudy (mathmale):

Please think about this and then try finding the zeros. If you'd like further help with this, please share your work with me, so that I could give you meaningful feedback.

OpenStudy (anonymous):